# Some new cases of Zilber–Pink in $Y(1)^{3}$

### v1

@@ Abstract.
We prove the Zilber–Pink conjecture for curves in $Y(1)^{3}$ that intersect a modular curve in the boundary. We also give an unconditional result for points having few places of supersingular reduction. Both results are proved using the G-function method for unlikely intersections.
@@ 1. Introduction
The Zilber–Pink conjecture is one of the central problems in arithmetic geometry today. It can be formulated in many different settings but, for (pure) Shimura varieties, we can state it as follows.
Conjecture 1 (Zilber–Pink).
@@
    $X_{i}=x$, for some $x$ in the set $\Sigma_{\rm CM}$ of singular moduli;
@@
where $\Phi_{N}(X,Y)\in\mathbb{Z}[X,Y]$ denotes the $N$th modular polynomial, defined by the property that two elliptic curves with $j$-invariants $j_{1}$ and $j_{2}$ are isogenous, with minimal isogeny of degree $N$, if and only if $\Phi_{N}(j_{1},j_{2})=0$. Special subvarieties defined by conditions only of type (2)—namely, those with no fixed coordinate—are examples of so-called strongly special subvarieties [CU05].
It follows, then, that the problem of Zilber–Pink in $Y(1)^{3}$ is reduced to the following, rather elegant conjecture.
Conjecture 2.
@@
Note that, by the preceding discussion, it suffices to consider only those points $s=(x_{1},x_{2},x_{3})\in Y(1)^{3}$ whose coordinates are non-singular moduli. For such points, the $M,N\in\mathbb{N}$ for which
$\Phi_{M}(x_{1},x_{2})=\Phi_{N}(x_{2},x_{3})=0$
are uniquely determined. We henceforth refer to such a point as modular and define
@@
By an ingenious argument, Habegger–Pila show that Conjecture 2 holds when $C$ is asymmetric, which (in this case) is to say that the degrees of the coordinate projections $C\to Y(1)$ are not all equal [HP12, Thm. 1]. In some sense, this covers most curves. On the other hand, it is not clear to us how to generalize their argument, or how to deal with symmetric curves specifically.
@@
With this in mind, the present authors have more recently explored an alternative approach, based on techniques of André and Bombieri using G-functions. In [DO22], the first and second authors showed that Conjecture 2 holds for any $C$ whose Zariski closure $\overline{C}$ in $X(1)^{3}\cong(\mathbb{P}^{1})^{3}$ (the Baily–Borel compactification) contains $(\infty,\infty,\infty)$. In [Pap], the third author observed that this could be extended to any curve $C$ for which $\overline{C}\setminus C$ contains a point $(x_{1},x_{2},x_{3})$ with $x_{1},x_{2},x_{3}\in\{\infty\}\cup\Sigma_{\rm CM}$. In other words, $\overline{C}$ intersects a special point in the boundary.
@@
Note that the boundary can naturally be identified with the disjoint union
$Y(1)^{2}\sqcup Y(1)^{2}\sqcup Y(1)^{2}\sqcup Y(1)\sqcup Y(1)\sqcup Y(1)\sqcup\{\infty\}$
@@
of products of modular curves and the point $\infty:=(\infty,\infty,\infty)$. Since $\overline{C}$ must always intersect the boundary, it is interesting to ask which intersection conditions facilitate our method. A natural next step in this direction is the case of a curve intersecting a special curve in the boundary.
@@
To make this precise, we say that a point in the boundary is modular if it belongs to one of the copies of $Y(1)^{2}$ and its coordinates $x_{1},x_{2}$ satisfy $\Phi_{N}(x_{1},x_{2})=0$ for some $N\in\mathbb{N}$. In this terminology, we prove the following.
Theorem 3.
@@
Conjecture 2 holds under the assumption that $\overline{C}\setminus C$ contains a modular point.
@@
However, our methods also yield a Zilber–Pink–type result without conditions at the boundary. In order to state it, let $C$ be as in the statement of Conjecture 2. Observe that we can always assume that $C$ contains a modular point $(b_{1},b_{2},b_{3})$. (Otherwise, there is nothing to prove!) Observe also that any modular point on $C$ belongs to $C(\overline{\mathbb{Q}})$. Thus, we denote by $K_{0}$ the number field $\mathbb{Q}(b_{1},b_{2},b_{3})$, and we say that a finite place $v$ of any finite extension of $K_{0}$ is a place of supersingular reduction if the elliptic curve with $j$-invariant $b_{1}$ (equivalently, $b_{2}$ or $b_{3}$) has supersingular reduction at $v$ (for the minimal Weierstrass model over $K_{v}$). For any such place $v$, lying over rational prime $p$, we denote by $|\cdot|_{v}$ the absolute value extending the standard $p$-adic absolute value on $\mathbb{Q}$. For any $\delta>0$, we say that a modular point $s=(x_{1},x_{2},x_{3})\in C$ has supersingular exponent $\delta$ if the set of places $v$ of $K_{0}(x_{1},x_{2},x_{3})$ of supersingular reduction for which
$|x_{1}-b_{1}|_{v},|x_{2}-b_{2}|_{v},|x_{3}-b_{3}|_{v}<1$
has cardinality at most $\Delta(s)^{\delta}$.
Theorem 4.
@@
Let $C$ and $(b_{1},b_{2},b_{3})$ be as above. There exists $\delta>0$ such that $C$ contains only finitely many modular points of supersingular exponent $\delta$.
@@
To prove Theorems 3 and 4, we obtain the following heights bounds. The derivations of Theorems 3 and 4 from these height bounds is now standard—facilitated by the so-called Pila–Zannier strategy and Masser–Wüstholz-type isogeny estimates—but we will give the details briefly in Section 5. In order to state them, we denote by $h(x_{1},x_{2},x_{3})$ the maximum of the absolute logarithmic heights $h(x_{i})$, as defined, for example, in [HP12, §2.2].
Theorem 5.
@@
Let $C$ satisfy the conditions of Theorem 3. There exist constants $c_{1}$ and $c_{2}$ such that, for any modular point $(x_{1},x_{2},x_{3})\in C$,
@@
Theorem 6.
Let $C$ and $(b_{1},b_{2},b_{3})$ be as in Theorem 4 and let $\epsilon>0$. There exist constants $c_{3}$, $c_{4}$ and $\delta$ such that, for any point modular point $s=(x_{1},x_{2},x_{3})\in C$ of supersingular exponent $\delta$,
@@
We refer to [DO23, Section 1.D] for a detailed history of the G-function method for problems of unlikely intersections. As alluded to above, the results herein are a continuation of that method. The original motivation, and one of the key novelties, of this paper was to carry out this method in the interior of a Shimura variety, as opposed to at a point of degeneration at the boundary. Theorem 4, which is close in spirit to [And95, Théorème 1] (and whose proof is very much inspired by the proof of the latter) is our main result in this direction. This line of inquiry will be followed up in a series of forthcoming papers of the third author [Pap25a, Pap25b, Pap25c]
@@
In Section 2, we give standard constructions pertaining to the G-functions method. In Section 3, we explain how to construct $v$-adic relations between the G-functions associated with elliptic schemes having isogenous fibres at a fixed point. In Section 4, we prove Theorems 5 and 6. In Section 5, we sketch the (now standard) arguments for deducing Theorems 3 and 4 from Theorems 5 and 6, respectively.
@@ Acknowledgements
The first author is indebted to Gregorio Baldi, with whom he had several interesting discussions around the idea of intersections at the boundary. For part of this work, he was supported by a grant from the School of Mathematics at the Institute for Advanced Studies, Princeton.
The second author’s work on this project was supported by the Engineering and Physical Sciences Research Council (EP/Y020758/1).
The third author started work on this project while supported by Michael Temkin’s ERC Consolidator Grant BirNonArchGeom (770922). For the majority of the work, he received funding from the European Union (ERC, SharpOS, 101087910) and the ISRAEL SCIENCE FOUNDATION (grant No. 2067/23). Latterly, he was supported by the Minerva Research Foundation Member Fund, while in residence at the Institute for Advanced Study, for the academic year 2025-26.
The authors thank Jacob Tsimerman for kindly answering questions related to his work [BT25].
@@ 2.1. Matrices
For a square matrix $M$, we denote its trace by $\operatorname{tr}(M)$ and its determinant by $\det(M)$. We denote its adjugate by $M^{\rm adj}$ (which, if $M$ is invertible, is equal to $\det(M)M^{-1}$) and its transpose by $M^{T}$. For any ring $R$, we denote by $\mathrm{M}_{2}(R)$ the ring of $2\times 2$ matrices with coefficients in $R$.
@@ 2.2. Fields
  • $K_{v}:=$ the completion of $K$ with respect to $v$;
  • $\mathcal{O}_{K_{v}}:=$ the ring of integers in $K_{v}$;
  • $\mathfrak{p}_{v}:=$ the prime ideal of $\mathcal{O}_{K}$ corresponding to $v$;
  • $k_{v}:=$ the residue field $\mathcal{O}_{K}/\mathfrak{p}_{v}$;
  • $W(k_{v}):=$ the ring of Witt vectors over $k_{v}$.
@@ 2.3. Elliptic curves
For an $R$-scheme $X\to\operatorname{Spec}R$ and an $R$-algebra $R\to S$, we denote by $X_{S}$ the base-change $X\times_{\operatorname{Spec}R}\operatorname{Spec}S$. For an elliptic curve $E$ over $K$, we say that $E$ has good reduction at $v\in\Sigma_{K,f}$ if the special fibre $\mathcal{E}_{k_{v}}$ of the minimal Weierstrass model $\mathcal{E}$ of $E_{K_{v}}$ over $\mathcal{O}_{K_{v}}$ is non-singular. We say $E$ has bad reduction at $v$ otherwise. If $\mathcal{E}_{k_{v}}$ is non-singular, it is an elliptic curve over $k_{v}$, and we say that $E$ has ordinary (resp. supersingular) reduction at $v$ if $\mathcal{E}_{k_{v}}$ is ordinary (resp. supersingular). All of these properties are preserved by isogenies over $K$.

### v2

@@ Abstract.
We prove the Zilber–Pink conjecture for curves in $Y(1)^{3}$ that intersect a modular curve in the boundary. We also give an unconditional result for unlikely intersection points having few places of supersingular reduction where they are close to a fixed base point. Both results are proved using the G-functions method for unlikely intersections.
@@ 1. Introduction
The aim of this paper is to prove new cases of the Zilber–Pink conjecture for $Y(1)^{3}$. The statement of this conjecture is that a geometrically irreducible algebraic curve $C\subset Y(1)^{3}\cong\mathbb{A}^{3}$, not contained in a proper special subvariety, has at most finitely many intersection points with special curves in $Y(1)^{3}$ (see below for the definition of special subvarieties in $Y(1)^{n}$). Habegger and Pila [HP12] have proved this conjecture for so-called asymmetric curves. More recently, the first- and second-named authors [DO25] and the third-named author [Pap] have proved it under restrictions on how the closure of $C$ in $(\mathbb{P}^{1})^{3}$ intersects the boundary $(\mathbb{P}^{1})^{3}\setminus\mathbb{A}^{3}$.
In this paper, we prove the conjecture when $C$ intersects a strongly special curve in the boundary (Theorem 1.3), in contrast to our previous results which require $C$ to intersect certain zero-dimensional subsets of the boundary. We also prove a version of the conjecture involving no condition at the boundary, but which only gives finiteness for points satisfying a condition on their places of supersingular reduction (Theorem 1.6).
The Zilber–Pink conjecture is one of the central problems in arithmetic geometry today. It can be formulated in many different settings. For (pure) Shimura varieties, it may be stated as follows.
Conjecture 1.1 (Zilber–Pink).
@@
    $X_{i}=x$, for some $x$ in the set $\Sigma_{\rm CM}$ of singular moduli (i.e. $j$-invariants of CM elliptic curves);
@@
where $\Phi_{N}(X,Y)\in\mathbb{Z}[X,Y]$ denotes the $N$th modular polynomial. This is defined by the property that two elliptic curves with $j$-invariants $j_{1}$ and $j_{2}$ are isogenous, with minimal isogeny of degree $N$, if and only if $\Phi_{N}(j_{1},j_{2})=0$. Special subvarieties defined by conditions only of type (2)—namely, those with no fixed coordinate—are examples of so-called strongly special subvarieties [CU05].
It follows that the problem of Zilber–Pink in $Y(1)^{3}$ is reduced to the following, rather elegant conjecture.
Conjecture 1.2.
@@
Because the André–Oort conjecture is known, it suffices to consider only those points $s=(x_{1},x_{2},x_{3})\in Y(1)^{3}$ whose coordinates are non-singular moduli. For such points, the positive integers $M,N\in\mathbb{N}$ for which
(1) $\Phi_{M}(x_{1},x_{2})=\Phi_{N}(x_{2},x_{3})=0$
are uniquely determined.
We henceforth refer to a point $s=(x_{1},x_{2},x_{3})\in(Y(1)\setminus\Sigma_{\rm CM})^{3}$ satisfying (1) as modular. For a modular point, we define
@@
By an ingenious argument, Habegger and Pila showed that Conjecture 1.2 holds when $C$ is asymmetric, which (in this case) is to say that the degrees of the coordinate projections $C\to Y(1)$ are not all equal [HP12, Thm. 1]. In some sense, this covers most curves. On the other hand, it is not clear to us how to generalize their argument, or how to deal with symmetric curves specifically.
@@
With this in mind, the present authors have more recently explored an alternative approach, based on techniques of André and Bombieri using G-functions. In [DO25], the first- and second-named authors showed that Conjecture 1.2 holds for any $C$ whose Zariski closure $\overline{C}$ in $X(1)^{3}\cong(\mathbb{P}^{1})^{3}$ (the Baily–Borel compactification) contains $(\infty,\infty,\infty)$. In [Pap], the third-named author observed that this could be extended to any curve $C$ for which $\overline{C}\setminus C$ contains a point $(x_{1},x_{2},x_{3})$ with $x_{1},x_{2},x_{3}\in\{\infty\}\cup\Sigma_{\rm CM}$. In other words, $\overline{C}$ intersects what we might call a special point in the boundary.
@@
Note that the boundary $X(1)^{3}\setminus Y(1)^{3}=(\mathbb{P}^{1})^{3}\setminus\mathbb{A}^{3}$ can naturally be identified with the disjoint union of three copies of $Y(1)^{2}$ (where exactly one coordinate is $\infty$), three copies of $Y(1)$ (where exactly two coordinates are $\infty$), and the singleton $\{(\infty,\infty,\infty)\}$.
@@
Since $\overline{C}$ must always intersect the boundary, it is interesting to ask which intersection conditions facilitate our method. A natural next step in this direction is the case of a curve intersecting a special curve in one of the copies of $Y(1)^{2}$ in the boundary.
@@
We say that a point in the boundary is modular if it belongs to one of the copies of $Y(1)^{2}$ and its coordinates $x_{1},x_{2}$ satisfy $\Phi_{N}(x_{1},x_{2})=0$ for some $N\in\mathbb{N}$. In this terminology, we prove the following.
Theorem 1.3.
@@
Conjecture 1.2 holds under the assumption that $\overline{C}\setminus C$ contains a modular point.
@@
Our methods also yield a Zilber–Pink-type result without conditions at the boundary, but with an extra restriction on the modular points of $C$ which we consider. In order to state this result, let $C$ be as in the statement of Conjecture 1.2. Observe that we can always assume that $C$ contains a modular point $s_{0}=(b_{1},b_{2},b_{3})$. (Otherwise, there is nothing to prove!) Observe also that any modular point on $C$ belongs to $C(\overline{\mathbb{Q}})$. We denote by $K_{0}$ the number field $\mathbb{Q}(b_{1},b_{2},b_{3})$, and we say that a finite place $v$ of any finite extension of $K_{0}$ is a place of supersingular reduction if the elliptic curve with $j$-invariant $b_{1}$ (equivalently, $b_{2}$ or $b_{3}$) has supersingular reduction at $v$ (for the minimal Weierstrass model over $K_{v}$). For any such place $v$, lying over a rational prime $p$, we denote by $|\cdot|_{v}$ the absolute value extending the standard $p$-adic absolute value on $\mathbb{Q}$.
Definition.
Given $\delta>0$, we say that a modular point $s=(x_{1},x_{2},x_{3})\in C$ has supersingular exponent $\delta$ if the set of places $v$ of $K_{0}(x_{1},x_{2},x_{3})$ of supersingular reduction for which the inequalities
(2) $|x_{1}-b_{1}|_{v},|x_{2}-b_{2}|_{v},|x_{3}-b_{3}|_{v}<1$
hold has cardinality at most $\Delta(s)^{\delta}$.
Remark 1.4.
This definition may be read as saying that $s$ is $v$-adically close to $s_{0}$ for only a small number of places of supersingular reduction. Note that, if $v$ is a place of supersingular reduction satisfying (2), then the elliptic curves with $j$-invariants $x_{1}$, $x_{2}$, $x_{3}$ also have supersingular reduction at $v$.
Remark 1.5.
By a result of Serre, the set of places of supersingular reduction has density zero. On the other hand, it is conjectured to be infinite, and this is known if $K_{0}$ has at least one real embedding [Elk89] (see [Cha18] for a related result). However, for any $K_{0}$ and any $s\neq s_{0}$, there are only finitely many places satisfying (2).
In this terminology, we prove the following.
Theorem 1.6.
@@
Let $C$ be as in Conjecture 1.2. Let $(b_{1},b_{2},b_{3})$ be a modular point in $C$. There exists $\delta>0$ such that $C$ contains only finitely many modular points of supersingular exponent $\delta$, with respect to $(b_{1},b_{2},b_{3})$.
The constant $\delta$ afforded to us by Theorem 1.6 is relatively explicit. Indeed, any $\delta$ satisfying the conditions of Theorem 5.2 is allowed. By Remark 4.3, therefore, we are allowed to take any $\delta<(72\ell)^{-1}$, where $\ell$ (depending on $C$) is defined in Section 4.2. It is straightforward to control $\ell$ in terms of the genus of $C$.
@@
To prove Theorems 1.3 and 1.6, we obtain the following height bounds. The derivations of Theorems 1.3 and 1.6 from these height bounds is now standard—using the so-called Pila–Zannier strategy and Masser–Wüstholz-type isogeny estimates—but we will give a few details in Section 5. In order to state our height bounds, we denote by $h(x_{1},x_{2},x_{3})$ the maximum of the absolute logarithmic heights of the $x_{i}$, as defined, for example, in [HP12, §2.2].
Theorem 1.7.
@@
Let $C$ satisfy the conditions of Theorem 1.3. There exist constants $c_{1}$ and $c_{2}$ such that, for any modular point $(x_{1},x_{2},x_{3})\in C$, we have
@@
Theorem 1.8.
Let $C$ and $(b_{1},b_{2},b_{3})$ be as in Theorem 1.6. There exist constants $c_{3}$, $c_{4}$ and $c_{5}$ such that, for any modular point $s=(x_{1},x_{2},x_{3})\in C$ of supersingular exponent $\delta$, we have
@@
As mentioned above, Theorems 1.7 and 1.8 are obtained via the so-called G-function method. We refer to [DO23, Section 1.D] for a detailed history of this method and its bearing on problems of unlikely intersections. The original motivation, and one of the key novelties, of this paper was to carry out the method around a base point in the interior of a Shimura variety, as opposed to a point of degeneration at the boundary. Theorem 1.6, which is close in spirit to [And95, Théorème 1] (and whose proof is very much inspired by the ideas introduced in [And89] and [And95]) is our main result in this direction. This line of inquiry is followed up in a series of recent papers of the third-named author [Pap25a, Pap25b, Pap25c].
Ideas of the proof
Consider a geometrically irreducible algebraic curve $C\subset Y(1)^{3}$ defined over a number field $K$, and a modular point $s_{0}$, either in $C(K)$ or in $\overline{C}(K)\setminus C(K)$. Let $\mathcal{E}_{1},\mathcal{E}_{2},\mathcal{E}_{3}$ denote families of elliptic curves over a suitable Zariski open subset of $C$ whose $j$-invariants are given by the three projections $C\to Y(1)$.
The Taylor expansions around $s_{0}$ of suitable linear combinations of complex periods of these familes of elliptic curves are G-functions (Section 2.6, following [And95, Sec. 1]). We need to interpret the evaluations of these G-functions at non-archimedean places in terms of periods which are suitably functorial, by choosing a suitable notion of “horizontal” sections of de Rham cohomology over an analytic neighbourhood $s_{0}$. At places of good reduction for $\mathcal{E}_{1,s_{0}}$, this is achieved using Berthelot and Ogus’s comparison between crystalline and de Rham cohomology [BO83, Ogu84]; in fact, the construction for non-archimedean places of good reduction closely parallels that for archimedean places, so that we are able to treat both in a uniform way in Section 3.2.1. At places of bad reduction, we instead use Tate uniformisation, similarly to the case of families with multiplicative degeneration at a $K$-point dealt with in [DO25, DO23] (Section 3.2.2).
Let $s\in C(L)$ be a modular point, for some finite extension $L$ of $K$. For each place $v$ of $L$, we use linear algebra calculations starting from the existence of an isogeny $\mathcal{E}_{1,s}\to\mathcal{E}_{2,s}$ to construct a non-trivial $v$-adic relation between the evaluations of our G-functions at $s$ (Sections 3.2.1 and 3.2.2). In order to obtain a global relation, we need to multiply these $v$-adic relations together. However, in general the relations at different places are different, and we do not know how many of them we will need to multiply together, so we do not have control of the degree of the global relation, and hence cannot prove Conjecture 1.2 in full.
This is where the extra conditions of Theorems 1.7 and 1.8 come in. In Theorem 1.7 (where $\mathcal{E}_{3}$ degenerates to $\mathbb{G}_{m}$ at $s_{0}$), a similar trick to that in [And89, Ch. X, 3.1 and 3.4] ensures that only archimedean places and places of bad reduction are relevant; the number of each of these is proportional to $L:\mathbb{Q}]$, so this gives the bound we need (Section 4.5, especially Section 4.5.5). In Theorem 1.8, by taking into account both isogenies $\mathcal{E}_{1,s}\to\mathcal{E}_{2,s}$ and $\mathcal{E}_{1,s}\to\mathcal{E}_{3,s}$, we are able to construct a single relation of fixed degree which is valid at all relevant places of ordinary good reduction. We are not able to do the same thing at places of supersingular good reduction, so instead we use the notion of supersingular exponent to impose a restriction on the number of places of supersingular reduction as a condition of the theorem (Sections 3.3 and 4.6).
@@
In Section 2, we give standard constructions pertaining to the G-function method. In Section 3, we explain how to construct $v$-adic relations between the G-functions associated with elliptic schemes having isogenous fibres at a fixed point. In Section 4, we prove Theorems 1.7 and 1.8. In Section 5, we sketch the (now standard) arguments for deducing Theorems 1.3 and 1.6 from Theorems 1.7 and 1.8, respectively. In the appendix, we prove some results on rigid analytic neighbourhoods to be used in this and future works. These statements are surely well-known, but we found it convenient to put them together in one place.
@@ Acknowledgements
The first-named author is indebted to Gregorio Baldi, with whom he had several interesting discussions around the idea of intersections at the boundary, and to Ananth Shankar, with whom he discussed several aspects of this work. He acknowledges the use of ChatGPT for conceptualisation. For part of this work, he was supported by a grant from the School of Mathematics at the Institute for Advanced Studies, Princeton.
The second-named author’s work on this project was supported by the Engineering and Physical Sciences Research Council (EP/Y020758/1).
The third-named author started work on this project while supported by Michael Temkin’s ERC Consolidator Grant BirNonArchGeom (770922). For the majority of the work, he received funding from the European Union (ERC, SharpOS, 101087910) and the ISRAEL SCIENCE FOUNDATION (grant No. 2067/23). Latterly, he was supported by the Minerva Research Foundation Member Fund, while in residence at the Institute for Advanced Study, for the academic year 2025-26.
The authors thank Jacob Tsimerman for kindly answering questions related to his work [BT25] and for several other enlightening discussions.
@@ 2.1. Matrices
For a square matrix $M$, we denote its trace by $\operatorname{tr}(M)$ and its determinant by $\det(M)$. We denote its adjugate by $M^{\rm adj}$ (which, if $M$ is invertible, is equal to $\det(M)M^{-1}$) and its transpose by $M^{T}$. For any ring $R$ and positive integer $n$, we denote by $\mathrm{M}_{n}(R)$ the ring of $n\times n$ matrices with coefficients in $R$.
@@ 2.2. Fields
  • $K_{v}:=$ the completion of $K$ with respect to $v$
  • $\mathcal{O}_{K_{v}}:=$ the ring of integers in $K_{v}$
  • $\mathfrak{p}_{v}:=$ the prime ideal of $\mathcal{O}_{K}$ corresponding to $v$
  • $k_{v}:=$ the residue field $\mathcal{O}_{K}/\mathfrak{p}_{v}$
  • $W(k_{v}):=$ the ring of Witt vectors over $k_{v}$
@@ 2.3. Elliptic curves
For a scheme $X\to\operatorname{Spec}R$ and an $R$-algebra $R\to S$, we denote by $X_{S}$ the base-change $X\times_{\operatorname{Spec}R}\operatorname{Spec}S$. For an elliptic curve $E$ over $K$, we say that $E$ has good reduction at $v\in\Sigma_{K,f}$ if the special fibre $\mathcal{E}_{k_{v}}$ of the minimal Weierstrass model $\mathcal{E}$ of $E_{K_{v}}$ over $\mathcal{O}_{K_{v}}$ is non-singular. We say $E$ has bad reduction at $v$ otherwise. If $\mathcal{E}_{k_{v}}$ is non-singular, it is an elliptic curve over $k_{v}$, and we say that $E$ has ordinary (resp. supersingular) reduction at $v$ if $\mathcal{E}_{k_{v}}$ is ordinary (resp. supersingular). All of these properties are preserved by isogenies over $K$.